I. Suppose population 1 has mean μί with variance σ2 and population 2 has mean μ2 with the same variance σ2. Let s and s denote the sample variances from two samples with size ni and n2 from the corresponding populations, respectively. Show that the pooled estimator 1i+(2-1)si pooled ni + n2 -2 is an unbiased estimator of σ2.
Answer #1

Given the population 1 with mean mu _{1} and population 2 with mean μ2 with the same variance sigma ^{2}.

Let X_{1},X_{2},..,X_{n1} and Y_{1},Y_{2},..,Y_{n2} be two random samples from two populations with common variance sigma ^{2}.

Let s_{1}^{2} and s_{2}^{2} denote the sample variances from two samples with size n_{1} and n_{2} from the corresponding populations.

TO PROVE:

The pooled estimator pool ed is an unbiased estimator of sigma ^{2}.

PROOF:

In order to prove that the pooled estimator is an unbiased estimator of sigma ^{2}, we first prove that the two variances are unbiased estimators of sigma ^{2}.

Let s_{1}^2=(1/(n_{1}-1))sum_{i=1}^{n_{1}}(X_{i}-X_{bar})^{2} be the variance of sample 1.

We know that (n_{1}-1)s_{1}^{2}/sigma ^{2}sim chi ^{2} distribution with 1 (n degrees of freedom.

Using the fact that the expected value of a chi ^{2} random variable with u degrees of freedom is u,

E[(n_{1}-1)s_{1}^{2}/sigma ^{2}]=n_{1}-1

Rightarrow ((n_{1}-1)/sigma ^{2})E(s_{1}^{2})=n_{1}-1

Rightarrow E(s_{1}^{2})=sigma ^{2}

Thus s_{1}^2 is an unbiased estimator of sigma ^{2}.   (1

In a similar manner,

Let s_{2}^2=(1/(n_{2}-1))sum_{i=1}^{n_{2}}(X_{i}-X_{bar})^{2} be the variance of sample 2.

We know that (n_{2}-1)s_{2}^{2}/sigma ^{2}sim chi ^{2} distribution with (n_{2}-1) degrees of freedom.

Using the fact that the expected value of a chi ^{2} random variable with u degrees of freedom is u,

  E[(n_{2}-1)s_{2}^{2}/sigma ^{2}]=n_{2}-1

Rightarrow ((n_{2}-1)/sigma ^{2})E(s_{2}^{2})=n_{2}-1

Rightarrow E(s_{2}^{2})=sigma ^{2}

Thus s_{2}^2 is an unbiased estimator of sigma ^{2}.   ightarrow (2)

Now let us consider the pooled variance pool ed

From (1), (2) and  linearity of expected value:

E[s_{pooled}^{2}]=E[[(n_{1}-1)s_{1}^{2}+(n_{2}-1)s_{2}^{2}]/(n_{1}+n_{2}-2)]

  =[(n_{1}-1)/(n_{1}+n_{2}-2)]E[s_{1}^{2}]+[(n_{2}-1)/(n_{1}+n_{2}-2)]E[s_{2}^{2}]

  =[(n_{1}-1)/(n_{1}+n_{2}-2)]sigma ^{2}+[(n_{2}-1)/(n_{1}+n_{2}-2)]sigma ^{2}

  =[[(n_{1}-1)+(n_{2}-1)]/(n_{1}+n_{2}-2)]sigma ^{2}

  =[(n_{1}+n_{2}-2)/(n_{1}+n_{2}-2)]sigma ^{2}

E[s_{pooled}^2]=sigma ^{2}

This shows that s_{pooled}^2 is an unbiased estimator of sigma ^{2}.